An agency reports that 11.2% of workers in a particular country belonged to unions. Suppose a sample of 400 workers is collected to determine whether union efforts to organize have increased union membership.

(a)Formulate the hypotheses that can be used to determine whether union membership has increased.

H0: p < 0.112

Ha: p ≥ 0.112

H0: p = 0.112

Ha: p ≠ 0.112

H0: p ≤ 0.112

Ha: p > 0.112

H0: p > 0.112

Ha: p ≤ 0.112

H0: p ≥ 0.112

Ha: p < 0.112

(b)If the sample results show that 48 of the workers belonged to unions, what is the p-value for your hypothesis test?

Find the value of the test statistic. (Round your answer to two decimal places.) ____________

Find the p-value. (Round your answer to four decimal places.) _____________

p-value =

(c)At

α = 0.05, what is your conclusion?

Reject H0. There is insufficient evidence to conclude that there has been an increase in union membership.

Do not reject H0. There is sufficient evidence to conclude that there has been an increase in union membership.    

Reject H0. There is sufficient evidence to conclude that there has been an increase in union membership.

Do not reject H0. There is insufficient evidence to conclude that there has been an increase in union membership.
Answer #1

Solution :

Given that,

= 0.112

1 - = 0.888

n = 400

x = 48

Level of significance = = 0.05

Point estimate = sample proportion = = x / n = 0.12

This a right (One) tailed test.

a)

The null and alternative hypothesis is,

H0: p ≤ 0.112

Ha: p > 0.112

b)

Test statistics

z = ( - ) / *(1-) / n

= ( 0.12 - 0.112) / (0.112*0.888) /400

= 0.51

P-value = P(Z>z)

= 1 - P(Z <z )

= 1- P(Z < 0.51)

= 1 - 0.6950

= 0.3050

c)

The p-value is p = 0.3050, and since p = 0.3050 > 0.05, it is concluded that the null hypothesis is not rejected.

Do not reject H0. There is insufficient evidence to conclude that there has been an increase in union membership.